A 300 kg beam is supported at each end. (The beam is horizontal). A 100 kg mass sits one third of the way from one end. Find the force acting upward at each end.
Answer #1

the reaction forces acting upwards are R1 and R2

balancing the force in vertical direction

R1 + R2 = 300*9.8 + 100*9.8

R1 + R2 = 3920 ---------(1)

TAKING MOment about the left end

100*9.8*l/3 + 300*9.8*l/2 = R2*l

R2 = 1796.7 N

R1 = 3920 - 1796. = 2123.3 N